OpenStudy (anonymous):

HELLLLLLLLLLP MEEEEEE=( At Joe's candy store, the total cost of 1 gumball and 1 lollipop is $0.74. The total cost of 1 chocolate bar and 1 lollipop is $0.92. The total cost of 1 gumball and 1 chocolate bar is $1.24. What is the cost in dollars of 1 chocolate bar?

OpenStudy (perl):

x + y = .74 z + y = .92 x + z = 1.24

OpenStudy (perl):

let x = price of gumball, y = price of lollipop , z= price of chocolate bar

OpenStudy (anonymous):

and then what?

OpenStudy (anonymous):

set the top equation equal to x and the second set it equal to z then plug the right side of both equations into the third equation.

OpenStudy (anonymous):

(.74-y)+(.92-y)=1.24

OpenStudy (anonymous):

Then solve for y

OpenStudy (anonymous):

And plug it into the other equations and solve for x and z